# RAD: Training an End-to-End Driving Policy via Large-Scale 3DGS-based Reinforcement Learning

### v1

@@ A.3 Metric Details
$ADD=\frac{1}{T_{safe}}\sum_{t=1}^{T_{safe}}d_{min}(t),$ (A11)
@@
$J_{long}=\frac{d^{2}v_{long}}{dt^{2}},$ (A12)
@@
A.4 More Qualitative Results
This section presents additional qualitative comparisons across various driving scenarios, including detours, crawling in dense traffic, traffic congestion, and U-turn maneuvers. The results highlight the effectiveness of our approach in generating smoother trajectories, enhancing collision avoidance, and improving adaptability in complex environments.
Figure A1: More Qualitative Results. Comparison between the IL-only policy and RAD in various driving scenarios: Detour (Rows 1-2), Crawl in Dense Traffic (Rows 3-4), Traffic Congestion (Rows 5-6), and U-turn(Rows 7-8).

### v2

@@ A.6 Metric Details
$ADD=\frac{1}{T_{safe}}\sum_{t=1}^{T_{safe}}d_{min}(t),$ (30)
@@
$J_{long}=\frac{d^{2}v_{long}}{dt^{2}},$ (31)
@@
A.7 More Qualitative Results
Fig. 7 presents additional qualitative comparisons across various driving scenarios, including detours, crawling in dense traffic, traffic congestion, and U-turn maneuvers. The results highlight the effectiveness of our approach in generating smoother trajectories, enhancing collision avoidance, and improving adaptability in complex environments.
Figure 7: More Qualitative Results. Comparison between the IL-only policy and RAD in various driving scenarios: Detour (Rows 1-2), Crawl in Dense Traffic (Rows 3-4), Traffic Congestion (Rows 5-6), and U-turn(Rows 7-8).